OpenStudy (anonymous):

What is maximum value of f(x,y)=3x+3y. Using the following inequalities. X>=0. Y>=0. 3x+2y<=18 6x+7y<=42

OpenStudy (anonymous):

draw lines 3x+2y=18 and 6x+7y=42 \[x \ge0, y \ge0~ gives~ graph~ lies~infirst~quadrant.\]

OpenStudy (anonymous):

We have graphed the equation. And worked it out with the answer 18. That's not the right answer.

OpenStudy (anonymous):

3x+2y=18 x=0,y=9,y=0, x=6 6x+7y=42 x=0,y=6,y=0,x=7 

OpenStudy (anonymous):

i think at E,f=3*14/3+3*2=14+6=20 but you need to show at all four points.

OpenStudy (anonymous):

Yes finally! Thank you so much.

OpenStudy (anonymous):

yw